OpenStudy (darkigloo):

calculus- how can i tell if a series is an alternating series by looking at it?

satellite73 (satellite73):

\[(-1)^n\] is usually there

OpenStudy (darkigloo):

will there always be a negative number?

satellite73 (satellite73):

just minus one to the n or n plus one

satellite73 (satellite73):

but it could be something like \(\sum \frac{(-2)^n}{3^n}\) say

satellite73 (satellite73):

on the other hand, you can always factor out the \((-1)^n\) so it is usually written that way

OpenStudy (darkigloo):

what if theres no exponent n in the numerator?

satellite73 (satellite73):

then no probably not you have an example?

satellite73 (satellite73):

alternating of course means it goes \[a_1-a_2+a_3-a_4+...\] easiest way to write it is \((-1)^na_n\)

OpenStudy (darkigloo):

i'll show you an example in 2 min when the problem comes back (online hw)

satellite73 (satellite73):

kk

OpenStudy (darkigloo):

cos(pi*k) / k^2 (-1)^(2k) / k^2 (-1)^k*cos(k)

satellite73 (satellite73):

ok i see

satellite73 (satellite73):

some kind of trick

OpenStudy (darkigloo):

sin(k) k / (-2)^k (-1)^k / ln(k)

satellite73 (satellite73):

first off \((-1)^{2k}=1\) since \(2k\) is even and minus one to an even power is 1 so that one does not alternate

OpenStudy (darkigloo):

ok

satellite73 (satellite73):

i have never seen a question like this i guess the question is "which ones alternate" right?

OpenStudy (darkigloo):

yes, mark each alternating series

satellite73 (satellite73):

cos(pi*k) / k^2 since \(\cos(\pi)=-1,\cos(2\pi)=1,\cos(3\pi)=-1\) etc, this is a tricky way to write an alternating series

satellite73 (satellite73):

(-1)^k*cos(k) this one has \((-1)^k\) so it alternates for sure

satellite73 (satellite73):

sin(k) does not alternate it is just sine

satellite73 (satellite73):

and the last two do alternate, since they have either \((-1)^k\) or \((-2)^k\)

OpenStudy (darkigloo):

thank you.

OpenStudy (darkigloo):

so if there is (-#)^k its always alternating ?

satellite73 (satellite73):

very odd question of course you could just check the first few terms to see yea \((-\xi)^k\) alternates

satellite73 (satellite73):

pretty clear why right? raise a negative number to an even power, it is positive, raise to an odd number, it is negative

OpenStudy (darkigloo):

ok thank you

satellite73 (satellite73):

yw

OpenStudy (zarkon):

i would be careful with \((-1)^k\cos(k)\) since the cosine can be positive and negative also